in 15 years dorothy will be 3 times as old as stan. 5 years ago the difference in their ages was 50 how old are dorothy and stan?
Answer by lwsshak3(11628)

in 15 years dorothy will be 3 times as old as stan. 5 years ago the difference in their ages was 50 how old are dorothy and stan?
**
let x=dorothy's present age
let y=stan's present age
..
15 years hence:
(x+15)=3(y+15)
x+15=3y+45
..
5 years ago:
(x-5)-(y-5)=50
x-5-y+5=50
x-y=50
x=50+y
..
x+15=3y+45
sub (50+y) for x
50+y+15=3y+45
2y=20
y=10
x=50+y=60
ans:
dorothy's present age=60
let y=stan's present age=10